OpenStudy (anonymous):

I could use some help on this question. @jim_thompson5910 @ganeshie8

OpenStudy (anonymous):

every angle of this hexagon is 120 degree.because \[\frac{ (6-2)*180 }{ 6 }=120\] 2x+120=180 2x=60 x=30